OpenStudy (anonymous):

log base 2 of 1/256 = ?

TheSmartOne (thesmartone):

\[\log _{2}\frac{ 1 }{ 256 }\] Is that your question

TheSmartOne (thesmartone):

@lawls

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i actually got the answer to that one. do you know how to do log base 27 of 81?

TheSmartOne (thesmartone):

Yeah sure lets see

TheSmartOne (thesmartone):

\[\log _{27}81=x\] \[27^x=81\] \[3^{3x}=3^4\] \[3x=4\] \[x=\frac{ 4 }{ 3 }\]

TheSmartOne (thesmartone):

@lawls